OpenStudy (anonymous):

The Computer Club spent $2,565 on mouse pads. The members plan to sell the mouse pads during the book fair. If they charge $9.50 for each mouse pad, how many must they sell in order to make a profit?

OpenStudy (anonymous):

9.5x = 2565 x = 2565/9.5 x = 270 They have to sell 270 to break even. Therefore in order to make a profit, they must sell 271 or higher. check... 9.5 * 271 = 2574.50 (not much of a profit, but 271 is where the profit making starts.